### v1

# Low-Rank Graph Contrastive Learning for Node Classification

Abstract
Graph Neural Networks (GNNs) have been widely used to learn node representations and with outstanding performance on various tasks such as node classification. However, noise, which inevitably exists in real-world graph data, would considerably degrade the performance of GNNs revealed by recent studies. In this work, we propose a novel and robust GNN encoder, Low-Rank Graph Contrastive Learning (LR-GCL). Our method performs transductive node classification in two steps. First, a low-rank GCL encoder named LR-GCL is trained by prototypical contrastive learning with low-rank regularization. Next, using the features produced by LR-GCL, a linear transductive classification algorithm is used to classify the unlabeled nodes in the graph. Our LR-GCL is inspired by the low frequency property of the graph data and its labels, and it is also theoretically motivated by our sharp generalization bound for transductive learning. To the best of our knowledge, our theoretical result is among the first to theoretically demonstrate the advantage of low-rank learning in graph contrastive learning supported by strong empirical performance. Extensive experiments on public benchmarks demonstrate the superior performance of LR-GCL and the robustness of the learned node representations. The code of LR-GCL is available at https://anonymous.4open.science/r/Low-Rank_Graph_Contrastive_Learning-64A6/.
1 Introduction
Graph Neural Networks (GNNs) have become popular tools for node representation learning in recent years [Kipf and Welling, 2017, Bruna et al., 2014, Hamilton et al., 2017, Xu et al., 2019]. Most prevailing GNNs  [Kipf and Welling, 2017, Zhu and Koniusz, 2020] leverage the graph structure and obtain the representation of nodes in a graph by utilizing the features of their connected nodes. Benefiting from such propagation mechanism, node representations obtained by GNN encoders have demonstrated superior performance on various downstream tasks such as semi-supervised node classification and node clustering.
Although GNNs have achieved great success in node representation learning, many existing GNN approaches do not consider the noise in the input graph. In fact, noise inherently exists in the graph data for many real-world applications. Such noise may be present in node attributes or node labels, which forms two types of noise, attribute noise and label noise. Recent works, such as [Patrini et al., 2017], have evidenced that noisy inputs hurt the generalization capability of neural networks. Moreover, noise in a subset of the graph data can easily propagate through the graph topology to corrupt the remaining nodes in the graph data. Nodes that are corrupted by noise or falsely labeled would adversely affect the representation learning of themselves and their neighbors.
While manual data cleaning and labeling could be remedies to the consequence of noise, they are expensive processes and difficult to scale, thus not able to handle almost infinite amount of noisy data online. Therefore, it is crucial to design a robust GNN encoder that could make use of noisy training data while circumventing the adverse effect of noise. In this paper, we propose a novel GCL encoder termed Low-Rank Graph Contrastive Learning (LR-GCL) to improve the robustness and the generalization capabilities of node representations for GNNs.
1.1 Contributions
Our contributions are as follows.
First, we present a novel and provable GCL encoder termed Low-Rank Graph Contrastive Learning (LR-GCL). Our algorithm is inspired by the low frequency property illustrated in Figure 1. That is, the low-rank projection of the ground truth clean labels possesses the majority of the information of the clean labels, and projection of the label noise is mostly uniform over all the eigenvectors of a kernel matrix used in classification. Inspired by this observation, LR-GCL adds the truncated nuclear norm as a low-rank regularization term in the loss function of the regular prototypical graph contrastive learning. As a result, the features produced by LR-GCL tend to be low-rank, and such low-rank features are the input to the linear transductive classification algorithm. We provide a novel generalization bound for the test loss on the unlabeled data, and our bound is among the first few works which exhibit the advantage of learning with low-rank features for transductive classification with the presence of noise.
Second, we provide strong theoretical guarantee on the generalization capability of the linear transductive algorithm with the low-rank features produced by LR-GCL as the input. Extensive experimental results on popular graph datasets evidence the advantage of LR-GCL over competing methods for node classification on noisy graph data.
Figure 1: Eigen-projection (first row) and signal concentration ratio (second row) on Cora, Citeseer, and Pubmed. To compute the eigen-projection, we first calculate the eigenvectors $\mathbf{U}$ of the feature gram matrix $\mathbf{K}=\mathbf{H}\mathbf{H}^{\top}$, then the eigen-projection value is computed by $\mathbf{p}=\frac{1}{C}\sum_{c=1}^{C}{\left\|{\mathbf{U}}^{\top}\tilde{}\mathbf{Y}^{(c)}\right\|}_{2}^{2}/{\left\|\tilde{}\mathbf{Y}^{(c)}\right\|}_{2}^{2}\in\mathbb{R}^{N}$, where $C$ is the number of classes, and ~$\tilde{}\mathbf{Y}\in\{0,1\}^{N\times C}$ is the one-hot clean labels of all the nodes, ~$\tilde{}\mathbf{Y}^{(c)}$ is the $c$-th column of ~$\tilde{}\mathbf{Y}$. With the presence of label noise $\mathbf{N}\in\mathbb{R}^{N\times C}$, the observed label matrix is $\mathbf{Y}=\tilde{}\mathbf{Y}+\mathbf{N}$. The eigen-projection $\mathbf{p}_{r}$ for $r\in[N]$ reflects the amount of the signal projected onto the $r$-th eigenvector of $\mathbf{K}$, and the signal concentration ratio of a rank $r$ reflects the proportion of signal projected onto the top $r$ eigenvectors of $\mathbf{K}$. The signal concentration ratio for rank $r$ is computed by ${\left\|\mathbf{p}^{(1:r)}\right\|}_{2}$, where $\mathbf{p}^{(1:r)}$ contains the first $r$ elements of $\mathbf{p}$. It is observed from the red curves in the first row that the projection of the ground truth clean labels mostly concentrates on the top eigenvectors of $\mathbf{K}$. On the other hand, the projection of label noise, computed by $\frac{1}{C}\sum_{c=1}^{C}{\left\|{\mathbf{U}}^{\top}\mathbf{N}^{(c)}\right\|}_{2}^{2}/{\left\|\mathbf{Y}^{(c)}\right\|}_{2}^{2}\in\mathbb{R}^{N}$, is relatively uniform over all the eigenvectors, as illustrated by the blue curves in the first row. For example, by the rank $r=0.2\min\left\{N,d\right\}$, the signal concentration ratio of ~$\tilde{}\mathbf{Y}$ for Cora, Citeseer, and Pubmed are $0.844$, $0.809$, and $0.784$ respectively. We refer to such property as the low frequency property, which suggests that we can learn a low-rank portion of the observed label $\mathbf{Y}$ which covers most information in the ground truth clean label while only learning a small portion of the label noise. Figure 3 in the supplementary further illustrates the low frequency property on more datasets.
2 Related Works
2.1 Graph Neural Networks
Graph neural networks (GNNs) have recently become popular tools for node representation learning. Given the difference in the convolution domain, current GNNs fall into two classes. The first class features spectral convolution [Bruna et al., 2014, Kipf and Welling, 2017], and the second class [Hamilton et al., 2017, Veličković et al., 2017, Xu et al., 2019] generates node representations by sampling and propagating features from their neighborhood. To learn node representation without node labels, contrastive learning has recently been applied to the training of GNNs [Suresh et al., 2021, Thakoor et al., 2021, Wang et al., 2022, Lee et al., 2022, Feng et al., 2022a, Zhang et al., 2023, Lin et al., 2023]. Most proposed graph contrastive learning methods [Veličković et al., 2019, Sun et al., 2019, Hu et al., 2019, Jiao et al., 2020, Peng et al., 2020, You et al., 2021, Jin et al., 2021, Mo et al., 2022] create multiple views of the unlabeled input graph and maximize agreement between the node representations of these views. For example, SFA [Zhang et al., 2023] manipulates the spectrum of the node embeddings to construct augmented views in graph contrastive learning. In addition to constructing node-wise augmented views, recent works [Xu et al., 2021, Guo et al., 2022, Li et al., 2021] propose to perform contrastive learning between node representations and semantic prototype representations [Snell et al., 2017, Arik and Pfister, 2020, Allen et al., 2019, Xu et al., 2020] to encode the global semantics information.
However, as pointed out by [Dai et al., 2021], the performance of GNNs can be easily degraded by noisy training data [NT et al., 2019]. Moreover, the adverse effects of noise in a subset of nodes can be exaggerated by being propagated to the remaining nodes through the network structure, exacerbating the negative impact of noise. Unlike previous GCL methods, we propose using contrastive learning to train GNN encoders that are robust to noise existing in the labels and attributes of nodes.
2.2 Existing Methods Handing Noisy Data
Previous works [Zhang et al., 2021] have shown that deep neural networks usually generalize badly when trained on input with noise. Existing literature on robust learning mostly fall into two categories. The first category [Patrini et al., 2017, Goldberger and Ben-Reuven, 2016] mitigates the effects of noisy inputs by correcting the computation of loss function, known as loss corruption. The second category aims to select clean samples from noisy inputs for the training [Malach and Shalev-Shwartz, 2017, Jiang et al., 2018, Yu et al., 2019, Li et al., 2020, Han et al., 2018], known as sample selection. To improve the performance of GNNs on graph data with noise, NRGNN[Dai et al., 2021] first introduces a graph edge predictor to predict missing links for connecting unlabeled nodes with labeled nodes. RTGNN [Qian et al., 2022] trains a robust GNN classifier with scarce and noisy node labels. It first classifies labeled nodes into clean and noisy ones and adopts reinforcement supervision to correct noisy labels. To improve the robustness of the node classifier on the dynamic graph, GraphSS [Zhuang and Al Hasan, 2022] proposes to generalize noisy supervision as a kind of self-supervised learning method, which regards the noisy labels, including both manual-annotated labels and auto-generated labels, as one kind of self-information for each node. Different from previous works, we aim to improve the robustness of GNN encoders for node classification by applying low-rank regularization during the training of the transductive classifier.
3 Problem Setup
3.1 Notations
An attributed graph consisting of $N$ nodes is formally represented by $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$, where $\mathcal{V}=\{v_{1},v_{2},\dots,v_{N}\}$ and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ denote the set of nodes and edges respectively. $\mathbf{X}\in\mathbb{R}^{N\times D}$ are the node attributes, and the attributes of each node is in $\mathbb{R}^{D}$. Let $\mathbf{A}\in\{0,1\}^{N\times N}$ be the adjacency matrix of graph $\mathcal{G}$, with $\mathbf{A}_{ij}=1$ if and only if $(v_{i},v_{j})\in\mathcal{E}$. $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ denotes the adjacency matrix for a graph with self-loops added. $\tilde{\mathbf{D}}$ denotes the diagonal degree matrix of $\tilde{\mathbf{A}}$. $[n]$ denotes all the natural numbers between $1$ and $N$ inclusively. ${\cal L}$ is a subset of $[N]$ of size $m$, and ${\cal U}$ is a subset of $[N]\setminus{\cal L}$ and $\left|{\cal U}\right|=u$. Let ${\cal V}_{{\cal L}}$ and ${\cal V}_{{\cal U}}$ denote the set of labeled nodes and unlabeled test nodes respectively, and $\left|{\cal V}_{{\cal L}}\right|=m$, $\left|{\cal V}_{{\cal U}}\right|=u$. Note that $m+u\leq N$, and it is not necessary that $m+u=N$ because there are usually validation nodes other than the labeled nodes and unlabeled test nodes. Let $\mathbf{u}\in\mathbb{R}^{N}$ be a vector, we use $\left[\mathbf{u}\right]_{{\cal A}}$ to denote a vector formed by elements of $\mathbf{u}$ with indices in ${\cal L}$ for ${\cal A}\subseteq[N]$ If $\mathbf{u}$ is a matrix, then $\left[\mathbf{u}\right]_{{\cal A}}$ denotes a submatrix formed by rows of $\mathbf{u}$ with row indices in ${\cal A}$. ${\left\|\cdot\right\|}_{\textup{F}}$ denotes the Frobenius norm of a matrix, and ${\left\|\cdot\right\|}_{p}$ denotes the $p$-norm of a vector.
3.2 Graph Convolution Network (GCN)
To learn the node representation from the attributes $\mathbf{X}$ and the graph structure $\mathbf{A}$, one simple yet effective neural network model is Graph Convolution Network (GCN). GCN is originally proposed for semi-supervised node classification, which consists of two graph convolution layers. In our work, we use GCN as the backbone of the proposed LR-GCL, which is the GCL encoder, to obtain node representation $\mathbf{H}\in\mathbb{R}^{N\times d}$, where the $i$-th row of $\mathbf{H}$ is the node representation of $v_{i}$. In this manner, the output of LR-GCL is $\mathbf{H}=g(\mathbf{X},\mathbf{A})=\sigma(\widehat{\mathbf{A}}\sigma(\widehat{\mathbf{A}}\mathbf{X}{\tilde{}\mathbf{W}}^{(0)}){\tilde{}\mathbf{W}}^{(1)})$, where $\widehat{\mathbf{A}}={\tilde{\mathbf{D}}}^{-1/2}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-1/2}$, ~${\tilde{}\mathbf{W}}^{(0)}$ and ~${\tilde{}\mathbf{W}}^{(1)}$ are the weight matrices, and $\sigma$ is the activation function ReLU. The robust and low-rank node representations produced by the LR-GCL are used to perform transductive node classification by a linear classifier. LR-GCL and the linear transductive node classification algorithm are detailed in Section 4.
3.3 Problem Description
Noise usually exists in the input node attributes or labels of real-world graphs, which degrades the quality of the node representation obtained by common GCL encoders and affects the performance of the classifier trained on such representations. We aim to obtain node representations robust to noise in two cases, where noise is present in either the labels of $\mathcal{V_{L}}$ or in the input node attributes $\mathbf{X}$. That is, we consider either noisy label or noisy input node attributes.
The goal of LR-GCL is to learn low-rank node representations by $\mathbf{H}=g(\mathbf{X},\mathbf{A})$ such that the node representations $\left\{\mathbf{h}_{i}\right\}_{i=1}^{N}$ are robust to noise in the above two cases, where $g(\cdot)$ is the LR-GCL encoder. In our work, $g$ is a two-layer GCN introduced in Section 3.2. The low-rank node representations by LR-GCL, $\mathbf{H}=\left\{\mathbf{h}_{1};\mathbf{h}_{2};\ldots;\mathbf{h}_{N}\right\}\in\mathbb{R}^{N\times d}$, are used for transductive node classification by a linear classifier. In transductive node classification, a linear transductive classifier is trained on $\mathcal{V_{L}}$, and then the classifier predicts the labels of the unlabeled test nodes in $\mathcal{V_{U}}$.
Figure 2: Illustration of the LR-GCL framework.
4 Methods
4.1 Low-Rank GCL: Low-Rank Graph Contrastive Learning
Preliminary of Prototypical GCL. The general node representation learning aims to train an encoder $g(\cdot)$, which is a two-layer Graph Convolution Neural Network (GCN) [Kipf and Welling, 2017], to generate discriminative node representations. In our work, we adopt contrastive learning to train the GCL encoder $g(\cdot)$. To perform contrastive learning, two different views, $G^{1}=(\mathbf{X}^{1},\mathbf{A}^{1})$ and $G^{2}=(\mathbf{X}^{2},\mathbf{A}^{2})$, are generated by node dropping, edge perturbation, and attribute masking. The representation of two generated views are denoted as $\mathbf{H}^{1}=g(\mathbf{X}^{1},\mathbf{A}^{1})$ and $\mathbf{H}^{2}=g(\mathbf{X}^{2},\mathbf{A}^{2})$, with $\mathbf{H}^{1}_{i}$ and $\mathbf{H}^{2}_{i}$ being the $i$-th row of $\mathbf{H}^{1}$ and $\mathbf{H}^{2}$, respectively. It is preferred that the mutual information between $\mathbf{H}^{1}$ and $\mathbf{H}^{2}$ is maximized. For computational reason, its lower bound is usually used as the objective for contrastive learning. We use InfoNCE [Li et al., 2021] as our node-wise contrastive loss. In addition to the node-wise contrastive learning, we also adopt prototypical contrastive learning [Li et al., 2021] to capture semantic information in the node representations, which is interpreted as maximizing the mutual information between node representation and a set of estimated cluster prototypes $\left\{\mathbf{c}_{1},...,\mathbf{c}_{{K}}\right\}$. Following [Li et al., 2021, Snell et al., 2017], we use $K$-means to cluster the node representations $\left\{\mathbf{h}_{i}\right\}_{i=1}^{n}$ into $K$ clusters and take the clustering centroid of the $k$-th cluster as the $k$-th prototype $\mathbf{c}_{k}=\frac{1}{|S_{k}|}\sum_{\mathbf{h}_{i}\in S_{k}}\mathbf{h}_{i}$ for all $k\in[K]$. The loss function of Prototypical GCL is comprised of two terms, $\mathcal{L}_{node}$, the loss function for node-wise contrastive learning, and $\mathcal{L}_{\textup{proto}}$, the prototypical contrastive learning loss, which are presented below:
$\small\mathcal{L}_{\textup{node}}$ $=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{s(\mathbf{H}^{1}_{i},\mathbf{H}^{2}_{i})}{s(\mathbf{H}^{1}_{i},\mathbf{H}^{2}_{i})+\sum_{j=1}^{N}s(\mathbf{H}^{1}_{i},\mathbf{H}^{2}_{j})},$
$\mathcal{L}_{\textup{proto}}$ $=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\mathbf{H}_{i}\cdot\mathbf{c}_{k}/\tau)}{\sum_{k=1}^{{K}}\exp(\mathbf{H}_{i}\cdot\mathbf{c}_{k}/\tau)}.$ (1)
Here $s(\mathbf{H}^{1}_{i},\mathbf{H}^{2}_{i})$ is the cosine similarity between two node representations, $\mathbf{H}^{1}_{i}$ and $\mathbf{H}^{2}_{i}$.
LR-GCL: Low-Rank Graph Contrastive Learning. LR-GCL aims to improve the robustness and generalization capability of the node representations of Prototypical GCL by enforcing the learned feature kernel to be low-rank. The gram matrix $\mathbf{K}$ of the node representations $\mathbf{H}\in\mathbb{R}^{N\times d}$ is calculated by $\mathbf{K}=\mathbf{H}^{\top}\mathbf{H}\in\mathbb{R}^{N\times N}$. Let $\left\{\widehat{\lambda}_{i}\right\}_{i=1}^{n}$ with $\widehat{\lambda}_{1}\geq\widehat{\lambda}_{2}\ldots\geq\widehat{\lambda}_{\min\left\{N,d\right\}}\geq\widehat{\lambda}_{\min\left\{N,d\right\}+1}=\ldots,=0$ be the eigenvalues of $\mathbf{K}$. In order to encourage the features $\mathbf{H}$ or the gram matrix $\mathbf{H}^{\top}\mathbf{H}$ to be low-rank, we explicitly add the truncated nuclear norm ${\left\|\mathbf{H}\right\|}_{r_{0}+1}\coloneqq\sum_{r=r_{0}}^{n}\widehat{\lambda}_{i}$ to the loss function of prototypical GCL. The starting rank $r_{0}<\min(n,d)$ is the rank of the features $\mathbf{H}$ we aim to obtain with the LR-GCL encoder, that is, if ${\left\|\mathbf{H}\right\|}_{r_{0}}=0$, then $\textup{rank}(\mathbf{H})=r_{0}$. Therefore, the overall loss function of LR-GCL is
$\mathcal{L}_{\textup{LR-GCL}}=\mathcal{L}_{\textup{node}}+\mathcal{L}_{\textup{proto}}+\tau{\left\|\mathbf{H}\right\|}_{r_{0}},$ (2)
where $\tau>0$ is the weighting parameter for the truncated nuclear norm ${\left\|\mathbf{H}\right\|}_{r_{0}}$. We summarize the training algorithm for the LR-GCL encoder in Algorithm 1. After finishing the training, we calculate the low-rank node feature by $\mathbf{H}^{(r_{0})}=g(\mathbf{A},\mathbf{X})$.
Motivation of Learning Low-Rank Features. Let ~$\tilde{}\mathbf{Y}\in\mathbb{R}^{N\times C}$ be the ground truth clean label matrix without noise. By the low frequency property illustrated in Figure 1, the projection of ~$\tilde{}\mathbf{Y}$ on the top $r$ eigenvectors of $\mathbf{K}$ with a small rank $r$, such as $r=0.2N$, covers the majority of the information in ~$\tilde{}\mathbf{Y}$. On the other hand, the projection of the label noise $\mathbf{N}$ are distributed mostly uniform across all the eigenvectors. This observation motivates low-rank features $\mathbf{H}$ or equivalently, the low-rank gram matrix $\mathbf{K}$. This is because the low-rank part of the feature matrix $\mathbf{H}$ or the gram matrix $\mathbf{K}$ covers the dominant information in the ground truth label ~$\tilde{}\mathbf{Y}$ while learning only a small portion of the label noise. Moreover, we remark that the regularization term ${\left\|\mathbf{H}\right\|}_{r_{0}}$ in the loss function (2) of LR-GCL is also theoretically motivated by the sharp upper bound for the test loss using a linear transductive classifier, presented as (A) in Theorem 4.1. A smaller ${\left\|\mathbf{H}\right\|}_{r_{0}}$ renders a smaller upper bound for the test loss, which ensures better generalization capability of the linear transductive classier to be introduced in the next subsection.
Algorithm 1 Low-Rank Graph Contrastive Learning (LR-GCL)
0:  The input attribute matrix $\mathbf{X}$, adjacency matrix $\mathbf{A}$, and the training epochs $t_{\text{max}}$.
0:  The parameters of LR-GCL encoder $g$
1:  Initialize the parameter of LR-GCL encoder $g$
2:  for $t\leftarrow 1$ to $t_{\text{max}}$ do
3:     Calculate node representations by $\mathbf{H}=g(\mathbf{X},\mathbf{A})$, generate augmented views $G^{1},G^{2}$, and calculate node representations $\mathbf{H}^{1}=g(\mathbf{X}^{1},\mathbf{A}^{1})$ and $\mathbf{H}^{2}=g(\mathbf{X}^{2},\mathbf{A}^{2})$
4:     Cluster node representations $\left\{\mathbf{h}_{i}\right\}_{i=1}^{n}$ into $K$ clusters $\left\{S_{k}\right\}_{k=1}^{K}$ with $K$-means clustering
5:     Update the prototype $\mathbf{c}_{k}$ as the centroid of $S_{k}$ by $\mathbf{c}_{k}=\frac{1}{|S_{k}|}\sum_{\mathbf{h}_{i}\in S_{k}}\mathbf{h}_{i}$ for all $k\in[K]$
6:     Calculate the eigenvalues $\left\{\lambda_{i}\right\}_{i=1}^{N}$ of the feature kernel $\mathbf{H}^{\top}\mathbf{H}$
7:     Update the parameters of LR-GCL encoder $g$ by one step of gradient descent on the loss $\mathcal{L}_{rep}$
8:  end for
9:  return The LR-GCL encoder $g$
4.2 Transductive Node Classification
In this section, we introduce a simple yet standard linear transductive node classification algorithm using the low-rank node representations $\mathbf{H}\in\mathbb{R}^{N\times d}$ produced by the LR-GCL encoder. We present strong theoretical result on the generalization bound for the test loss for our low-rank transductive algorithm with the presence of label noise.
We first give basic notations for our algorithm. Let $\mathbf{y}_{i}\in\mathbb{R}^{C}$ be the observed one-hot class label vector for node $v_{i}$ for all $i\in[N]$, and define $\mathbf{Y}\coloneqq\left[\mathbf{y}_{1};\mathbf{y}_{2};\ldots\mathbf{y}_{N}\right]\in\mathbb{R}^{N\times C}$ be the observed label matrix which may contain label noise $\mathbf{N}\in\mathbb{R}^{N\times C}$. We define $\mathbf{F}(\mathbf{W})=\mathbf{H}\mathbf{W}$ as the linear output of the transductive classier with $\mathbf{W}\in\mathbb{R}^{d\times C}$ being the weight matrix for the classifier. Our transductive classifier uses $\textup{softmax}(\mathbf{F}(\mathbf{W}))\in\mathbb{R}^{N\times C}$ for prediction of the labels of the test nodes. We train the transductive classifier by minimizing the regular cross-entropy on the labeled nodes via
$\min\limits_{\mathbf{W}}L(\mathbf{W})=\frac{1}{m}\sum_{v_{i}\in\mathcal{V_{L}}}\textup{KL}\left(\mathbf{y}_{i},\left[\textup{softmax}\left(\mathbf{H}\mathbf{W}\right)\right]_{i}\right),$ (3)
where KL is the KL divergence between the label $\mathbf{y}_{i}$ and the softmax of the classifier output at node $v_{i}$. We use a regular gradient descent to optimize (3) with a learning rate $\eta\in(0,\frac{1}{\widehat{\lambda}_{1}})$. We define a matrix ¯$\bar{}\mathbf{Y}(r_{0})\in\mathbb{R}^{N\times C}$ as the orthogonal projection of $\mathbf{Y}$ onto the top-$r_{0}$ eigenvectors of $\mathbf{K}$, that is, ¯$\bar{}\mathbf{Y}(r_{0})=\mathbf{U}^{(r_{0})}\left(\mathbf{U}^{(r_{0})}\right)^{\top}\mathbf{Y}$, where $\mathbf{U}^{(r_{0})}\in\mathbb{R}^{N\times r_{0}}$ is the matrix formed by the eigenvectors of $\mathbf{K}$ corresponding to the top-$r_{0}$ eigenvalues of $\mathbf{K}$. $\mathbf{W}$ is initialized by $\mathbf{W}^{(0)}=\mathbf{0}$, and at the $t$-th iteration of gradient descent for $t\geq 1$, $\mathbf{W}$ is updated by $\mathbf{W}^{(t)}=\mathbf{W}^{(t-1)}-\eta\nabla_{\mathbf{W}}L(\mathbf{W})|_{\mathbf{W}=\mathbf{W}^{(t-1)}}$.
Define $\mathbf{F}(\mathbf{W},t)\coloneqq\mathbf{H}\mathbf{W}^{(t)}$ as the output of the classifier after the $t$-th iteration of gradient descent for $t\geq 1$. We have the following theoretical result on the loss of the unlabeled test nodes ${\cal V}_{{\cal U}}$ measured by the gap between $\mathbf{F}(\mathbf{W},t)$ and ¯$\bar{}\mathbf{Y}(r_{0})$ when using the low-rank feature $\mathbf{H}$ with $r_{0}\in[n]$.
Theorem 4.1.
Let $m\geq cN$ for a constant $c\in(0,1)$, and $r_{0}\in[n]$. Assume that a set ${\cal L}$ with $\left|{\cal L}\right|=m$ is sampled uniformly without replacement from $[N]$, and a set ${\cal U}$ with $\left|{\cal U}\right|=u$ are sampled uniformly without replacement from $[N]\setminus{\cal L}$ and $m+u\leq N$. Then for every $x>0$, with probability at least $1-\exp(-x)$, after the $t$-th iteration of gradient descent for all $t\geq 1$, we have
${\cal U}_{\textup{test}}(t)\coloneqq\frac{1}{u}{\left\|\left[\mathbf{F}(\mathbf{W},t)-\mathbf{Y}\right]_{{\cal U}}\right\|}_{\textup{F}}^{2}$
$\leq\frac{c_{0}}{m}{\left\|\left(\mathbf{I}_{N}-\eta\mathbf{K}\right)^{t}\mathbf{Y}\right\|}_{\textup{F}}+c_{1}r_{0}\left(\frac{1}{u}+\frac{1}{m}\right)$
$+\left(\sqrt{\frac{{\left\|\mathbf{H}\right\|}_{r_{0}}}{u}}+\sqrt{\frac{{\left\|\mathbf{H}\right\|}_{r_{0}}}{m}}\right)+\frac{c_{2}x}{u},$ (4)
where $c_{0},c_{1},c_{2}$ are positive numbers depending on $\mathbf{U}$, $\left\{\widehat{\lambda}_{i}\right\}_{i=1}^{r_{0}}$, and $\tau_{0}$ with $\tau_{0}^{2}=\max_{i\in[N]}\mathbf{K}_{ii}$.
This theorem is proved in Section A of the supplementary. It is noted that $\frac{1}{u}{\left\|\left[\mathbf{F}(\mathbf{W},t)-\bar{}\mathbf{Y}(r_{0})\right]_{{\cal U}}\right\|}_{\textup{F}}^{2}$ is the test loss of the unlabeled nodes measured by the distance between the classifier output $\mathbf{F}(\mathbf{W},t)$ and $bY$. We remark that the truncated nuclear norm ${\left\|\mathbf{H}\right\|}_{r_{0}}$ appears on the RHS of the upper bound (A), theoretically justifying why we learn the low-rank features $\mathbf{H}$ of the LR-GCL by adding the ${\left\|\mathbf{H}\right\|}_{r_{0}}$ to the loss of our LR-GCL. Moreover, when the low frequency property holds, which is always the case as demonstrated by Figure 1 and Figure 3 in the supplementary, ${\left\|\left(\mathbf{I}_{N}-\eta\mathbf{K}\right)^{t}\mathbf{Y}\right\|}_{\textup{F}}$ would be very small with enough iteration number $t$.
In our empirical study in the next section, we search for the rank $r_{0}$ for the truncated nuclear norm by standard cross-validation for all the graph data sets. In Table 3 of our experimental results, it is observed that the best rank $r_{0}$ is always between $0.1\min\left\{N,d\right\}$ and $0.3\min\left\{N,d\right\}$. The overall framework of LR-GCL is illustrated in Figure 2.
5 Experiments
Table 1: Performance comparison for node classification on Cora, Citeseer, PubMed, and Wiki-CS with asymmetric label noise, symmetric label noise, and attribute noise.
Noise Level
0 40 60 80
Dataset Methods - Asymmetric Symmetric Attribute Asymmetric Symmetric Attribute Asymmetric Symmetric Attribute
Cora GCN 0.815$\pm$0.005 0.547$\pm$0.015 0.636$\pm$0.007 0.639$\pm$0.008 0.405$\pm$0.014 0.517$\pm$0.010 0.439$\pm$0.012 0.265$\pm$0.012 0.354$\pm$0.014 0.317$\pm$0.013
[…]
GCE 0.819$\pm$0.004 0.573$\pm$0.011 0.652$\pm$0.008 0.650$\pm$0.014 0.449$\pm$0.011 0.509$\pm$0.011 0.445$\pm$0.015 0.280$\pm$0.013 0.353$\pm$0.013 0.325$\pm$0.015
UnionNET 0.820$\pm$0.006 0.569$\pm$0.014 0.664$\pm$0.007 0.653$\pm$0.012 0.452$\pm$0.010 0.541$\pm$0.010 0.450$\pm$0.009 0.283$\pm$0.014 0.370$\pm$0.011 0.320$\pm$0.012
NRGNN 0.822$\pm$0.006 0.571$\pm$0.019 0.676$\pm$0.007 0.645$\pm$0.012 0.470$\pm$0.014 0.548$\pm$0.014 0.451$\pm$0.011 0.282$\pm$0.022 0.373$\pm$0.012 0.326$\pm$0.010
RTGNN 0.828$\pm$0.003 0.570$\pm$0.010 0.682$\pm$0.008 0.678$\pm$0.011 0.474$\pm$0.011 0.555$\pm$0.010 0.457$\pm$0.009 0.280$\pm$0.011 0.386$\pm$0.014 0.342$\pm$0.016
SUGRL 0.834$\pm$0.005 0.564$\pm$0.011 0.674$\pm$0.012 0.675$\pm$0.009 0.468$\pm$0.011 0.552$\pm$0.011 0.452$\pm$0.012 0.280$\pm$0.012 0.381$\pm$0.012 0.338$\pm$0.014
MERIT 0.831$\pm$0.005 0.560$\pm$0.008 0.670$\pm$0.008 0.671$\pm$0.009 0.467$\pm$0.013 0.547$\pm$0.013 0.450$\pm$0.014 0.277$\pm$0.013 0.385$\pm$0.013 0.335$\pm$0.009
ARIEL 0.843$\pm$0.004 0.573$\pm$0.013 0.681$\pm$0.010 0.675$\pm$0.009 0.471$\pm$0.012 0.553$\pm$0.012 0.455$\pm$0.014 0.284$\pm$0.014 0.389$\pm$0.013 0.343$\pm$0.013
SFA 0.839$\pm$0.010 0.564$\pm$0.011 0.677$\pm$0.013 0.676$\pm$0.015 0.473$\pm$0.014 0.549$\pm$0.014 0.457$\pm$0.014 0.282$\pm$0.016 0.389$\pm$0.013 0.344$\pm$0.017
Sel-Cl 0.828$\pm$0.002 0.570$\pm$0.010 0.685$\pm$0.012 0.676$\pm$0.009 0.472$\pm$0.013 0.554$\pm$0.014 0.455$\pm$0.011 0.282$\pm$0.017 0.389$\pm$0.013 0.341$\pm$0.015
Jo-SRC 0.825$\pm$0.005 0.571$\pm$0.006 0.684$\pm$0.013 0.679$\pm$0.007 0.473$\pm$0.011 0.556$\pm$0.008 0.458$\pm$0.012 0.285$\pm$0.013 0.387$\pm$0.018 0.345$\pm$0.018
GRAND+ 0.858$\pm$0.006 0.570$\pm$0.009 0.682$\pm$0.007 0.678$\pm$0.011 0.472$\pm$0.010 0.554$\pm$0.008 0.456$\pm$0.012 0.284$\pm$0.015 0.387$\pm$0.015 0.345$\pm$0.013
LR-GCL 0.858$\pm$0.006 0.589$\pm$0.011 0.713$\pm$0.007 0.695$\pm$0.011 0.492$\pm$0.011 0.587$\pm$0.013 0.477$\pm$0.012 0.306$\pm$0.012 0.419$\pm$0.012 0.363$\pm$0.011
Citeseer GCN 0.703$\pm$0.005 0.475$\pm$0.023 0.501$\pm$0.013 0.529$\pm$0.009 0.351$\pm$0.014 0.341$\pm$0.014 0.372$\pm$0.011 0.291$\pm$0.022 0.281$\pm$0.019 0.290$\pm$0.014
[…]
GCE 0.705$\pm$0.004 0.490$\pm$0.016 0.512$\pm$0.014 0.540$\pm$0.014 0.362$\pm$0.015 0.352$\pm$0.010 0.381$\pm$0.009 0.309$\pm$0.012 0.285$\pm$0.014 0.285$\pm$0.011
UnionNET 0.706$\pm$0.006 0.499$\pm$0.015 0.547$\pm$0.014 0.545$\pm$0.013 0.379$\pm$0.013 0.399$\pm$0.013 0.379$\pm$0.012 0.322$\pm$0.021 0.302$\pm$0.013 0.290$\pm$0.012
NRGNN 0.710$\pm$0.006 0.498$\pm$0.015 0.546$\pm$0.015 0.538$\pm$0.011 0.382$\pm$0.016 0.412$\pm$0.016 0.377$\pm$0.012 0.336$\pm$0.021 0.309$\pm$0.018 0.284$\pm$0.009
RTGNN 0.746$\pm$0.008 0.498$\pm$0.007 0.556$\pm$0.007 0.550$\pm$0.012 0.392$\pm$0.010 0.424$\pm$0.013 0.390$\pm$0.014 0.348$\pm$0.017 0.308$\pm$0.016 0.302$\pm$0.011
SUGRL 0.730$\pm$0.005 0.493$\pm$0.011 0.541$\pm$0.011 0.544$\pm$0.010 0.376$\pm$0.009 0.421$\pm$0.009 0.388$\pm$0.009 0.339$\pm$0.010 0.305$\pm$0.010 0.300$\pm$0.009
MERIT 0.740$\pm$0.007 0.496$\pm$0.012 0.536$\pm$0.012 0.542$\pm$0.010 0.383$\pm$0.011 0.425$\pm$0.011 0.387$\pm$0.008 0.344$\pm$0.014 0.301$\pm$0.014 0.295$\pm$0.009
SFA 0.740$\pm$0.011 0.502$\pm$0.014 0.532$\pm$0.015 0.547$\pm$0.013 0.390$\pm$0.014 0.433$\pm$0.014 0.389$\pm$0.012 0.347$\pm$0.016 0.312$\pm$0.015 0.299$\pm$0.013
ARIEL 0.727$\pm$0.007 0.500$\pm$0.008 0.550$\pm$0.013 0.548$\pm$0.008 0.391$\pm$0.009 0.427$\pm$0.012 0.389$\pm$0.014 0.349$\pm$0.014 0.307$\pm$0.013 0.299$\pm$0.013
Sel-Cl 0.725$\pm$0.008 0.499$\pm$0.012 0.551$\pm$0.010 0.549$\pm$0.008 0.389$\pm$0.011 0.426$\pm$0.008 0.391$\pm$0.020 0.350$\pm$0.018 0.310$\pm$0.015 0.300$\pm$0.017
Jo-SRC 0.730$\pm$0.005 0.500$\pm$0.013 0.555$\pm$0.011 0.551$\pm$0.011 0.394$\pm$0.013 0.425$\pm$0.013 0.393$\pm$0.013 0.351$\pm$0.013 0.305$\pm$0.018 0.303$\pm$0.013
GRAND+ 0.756$\pm$0.004 0.497$\pm$0.010 0.553$\pm$0.010 0.552$\pm$0.011 0.390$\pm$0.013 0.422$\pm$0.013 0.387$\pm$0.013 0.348$\pm$0.013 0.309$\pm$0.014 0.302$\pm$0.012
LR-GCL 0.757$\pm$0.010 0.520$\pm$0.013 0.581$\pm$0.013 0.570$\pm$0.007 0.410$\pm$0.014 0.455$\pm$0.014 0.406$\pm$0.012 0.369$\pm$0.012 0.335$\pm$0.014 0.318$\pm$0.010
PubMed GCN 0.790$\pm$0.007 0.584$\pm$0.022 0.574$\pm$0.012 0.595$\pm$0.012 0.405$\pm$0.025 0.386$\pm$0.011 0.488$\pm$0.013 0.305$\pm$0.022 0.295$\pm$0.013 0.423$\pm$0.013
[…]
GCE 0.792$\pm$0.009 0.589$\pm$0.018 0.581$\pm$0.011 0.590$\pm$0.014 0.430$\pm$0.012 0.399$\pm$0.012 0.491$\pm$0.010 0.311$\pm$0.021 0.301$\pm$0.011 0.424$\pm$0.012
UnionNET 0.793$\pm$0.008 0.603$\pm$0.020 0.620$\pm$0.012 0.592$\pm$0.012 0.445$\pm$0.022 0.424$\pm$0.013 0.489$\pm$0.015 0.313$\pm$0.025 0.327$\pm$0.015 0.435$\pm$0.009
NRGNN 0.797$\pm$0.008 0.602$\pm$0.022 0.618$\pm$0.013 0.603$\pm$0.008 0.443$\pm$0.012 0.434$\pm$0.012 0.499$\pm$0.009 0.330$\pm$0.023 0.325$\pm$0.013 0.433$\pm$0.011
RTGNN 0.797$\pm$0.004 0.610$\pm$0.008 0.622$\pm$0.010 0.614$\pm$0.012 0.455$\pm$0.010 0.455$\pm$0.011 0.501$\pm$0.011 0.335$\pm$0.013 0.338$\pm$0.017 0.452$\pm$0.013
SUGRL 0.819$\pm$0.005 0.603$\pm$0.013 0.615$\pm$0.013 0.615$\pm$0.010 0.445$\pm$0.011 0.441$\pm$0.011 0.501$\pm$0.007 0.321$\pm$0.009 0.321$\pm$0.009 0.446$\pm$0.010
MERIT 0.801$\pm$0.004 0.593$\pm$0.011 0.612$\pm$0.011 0.613$\pm$0.011 0.447$\pm$0.012 0.443$\pm$0.012 0.497$\pm$0.009 0.328$\pm$0.011 0.323$\pm$0.011 0.445$\pm$0.009
ARIEL 0.800$\pm$0.003 0.610$\pm$0.013 0.622$\pm$0.010 0.615$\pm$0.011 0.453$\pm$0.012 0.453$\pm$0.012 0.502$\pm$0.014 0.331$\pm$0.014 0.336$\pm$0.018 0.457$\pm$0.013
SFA 0.804$\pm$0.010 0.596$\pm$0.011 0.615$\pm$0.011 0.609$\pm$0.011 0.447$\pm$0.014 0.446$\pm$0.017 0.499$\pm$0.014 0.330$\pm$0.011 0.327$\pm$0.011 0.447$\pm$0.014
Sel-Cl 0.799$\pm$0.005 0.605$\pm$0.014 0.625$\pm$0.012 0.614$\pm$0.012 0.455$\pm$0.014 0.449$\pm$0.010 0.502$\pm$0.008 0.334$\pm$0.021 0.332$\pm$0.014 0.456$\pm$0.014
Jo-SRC 0.801$\pm$0.005 0.613$\pm$0.010 0.624$\pm$0.013 0.617$\pm$0.013 0.453$\pm$0.008 0.455$\pm$0.013 0.504$\pm$0.013 0.330$\pm$0.015 0.334$\pm$0.018 0.459$\pm$0.018
GRAND+ 0.845$\pm$0.006 0.610$\pm$0.011 0.624$\pm$0.013 0.617$\pm$0.013 0.453$\pm$0.008 0.453$\pm$0.011 0.503$\pm$0.010 0.331$\pm$0.014 0.337$\pm$0.013 0.458$\pm$0.014
LR-GCL 0.845$\pm$0.009 0.637$\pm$0.014 0.645$\pm$0.015 0.637$\pm$0.011 0.479$\pm$0.011 0.484$\pm$0.013 0.526$\pm$0.011 0.356$\pm$0.011 0.360$\pm$0.012 0.482$\pm$0.014
Coauthor-CS GCN 0.918$\pm$0.001 0.645$\pm$0.009 0.656$\pm$0.006 0.702$\pm$0.010 0.511$\pm$0.013 0.501$\pm$0.009 0.531$\pm$0.010 0.429$\pm$0.022 0.389$\pm$0.011 0.415$\pm$0.013
[…]
GCE 0.922$\pm$0.003 0.662$\pm$0.017 0.659$\pm$0.007 0.705$\pm$0.014 0.515$\pm$0.016 0.502$\pm$0.007 0.539$\pm$0.009 0.443$\pm$0.017 0.389$\pm$0.012 0.412$\pm$0.011
UnionNET 0.918$\pm$0.002 0.669$\pm$0.023 0.671$\pm$0.013 0.706$\pm$0.012 0.525$\pm$0.011 0.529$\pm$0.011 0.540$\pm$0.012 0.458$\pm$0.015 0.401$\pm$0.011 0.420$\pm$0.007
NRGNN 0.919$\pm$0.002 0.678$\pm$0.014 0.689$\pm$0.009 0.705$\pm$0.012 0.545$\pm$0.021 0.556$\pm$0.011 0.546$\pm$0.011 0.461$\pm$0.012 0.410$\pm$0.012 0.417$\pm$0.007
RTGNN 0.920$\pm$0.005 0.678$\pm$0.012 0.691$\pm$0.009 0.712$\pm$0.008 0.559$\pm$0.010 0.569$\pm$0.011 0.560$\pm$0.008 0.455$\pm$0.015 0.415$\pm$0.015 0.412$\pm$0.014
SUGRL 0.922$\pm$0.005 0.675$\pm$0.010 0.695$\pm$0.010 0.714$\pm$0.006 0.550$\pm$0.011 0.560$\pm$0.011 0.561$\pm$0.007 0.449$\pm$0.011 0.411$\pm$0.011 0.429$\pm$0.008
MERIT 0.924$\pm$0.004 0.679$\pm$0.011 0.689$\pm$0.008 0.709$\pm$0.005 0.552$\pm$0.014 0.562$\pm$0.014 0.562$\pm$0.011 0.452$\pm$0.013 0.403$\pm$0.013 0.426$\pm$0.005
ARIEL 0.925$\pm$0.004 0.682$\pm$0.011 0.699$\pm$0.009 0.712$\pm$0.005 0.555$\pm$0.011 0.566$\pm$0.011 0.556$\pm$0.011 0.454$\pm$0.014 0.415$\pm$0.019 0.427$\pm$0.013
SFA 0.925$\pm$0.009 0.682$\pm$0.011 0.690$\pm$0.012 0.715$\pm$0.012 0.555$\pm$0.015 0.567$\pm$0.014 0.565$\pm$0.013 0.458$\pm$0.013 0.402$\pm$0.013 0.429$\pm$0.015
Sel-Cl 0.922$\pm$0.008 0.684$\pm$0.009 0.694$\pm$0.012 0.714$\pm$0.010 0.557$\pm$0.013 0.568$\pm$0.013 0.566$\pm$0.010 0.457$\pm$0.013 0.412$\pm$0.017 0.425$\pm$0.009
Jo-SRC 0.921$\pm$0.005 0.684$\pm$0.011 0.695$\pm$0.004 0.709$\pm$0.007 0.560$\pm$0.011 0.566$\pm$0.011 0.561$\pm$0.009 0.456$\pm$0.013 0.410$\pm$0.018 0.428$\pm$0.010
GRAND+ 0.927$\pm$0.004 0.682$\pm$0.011 0.693$\pm$0.006 0.715$\pm$0.008 0.554$\pm$0.008 0.568$\pm$0.013 0.557$\pm$0.011 0.455$\pm$0.012 0.416$\pm$0.013 0.428$\pm$0.011
LR-GCL 0.933$\pm$0.006 0.699$\pm$0.015 0.721$\pm$0.011 0.742$\pm$0.015 0.575$\pm$0.014 0.595$\pm$0.018 0.588$\pm$0.015 0.469$\pm$0.015 0.438$\pm$0.015 0.453$\pm$0.017
ogbn-arxiv GCN 0.717$\pm$0.003 0.401$\pm$0.014 0.421$\pm$0.014 0.478$\pm$0.010 0.336$\pm$0.011 0.346$\pm$0.021 0.339$\pm$0.012 0.286$\pm$0.022 0.256$\pm$0.010 0.294$\pm$0.013
[…]
GCE 0.720$\pm$0.004 0.410$\pm$0.018 0.428$\pm$0.008 0.480$\pm$0.014 0.348$\pm$0.019 0.344$\pm$0.019 0.342$\pm$0.015 0.310$\pm$0.014 0.260$\pm$0.011 0.275$\pm$0.015
UnionNET 0.724$\pm$0.006 0.429$\pm$0.021 0.449$\pm$0.007 0.485$\pm$0.012 0.362$\pm$0.018 0.367$\pm$0.008 0.340$\pm$0.009 0.332$\pm$0.019 0.269$\pm$0.013 0.280$\pm$0.012
NRGNN 0.721$\pm$0.006 0.449$\pm$0.014 0.466$\pm$0.009 0.485$\pm$0.012 0.371$\pm$0.020 0.379$\pm$0.008 0.342$\pm$0.011 0.330$\pm$0.018 0.271$\pm$0.018 0.300$\pm$0.010
RTGNN 0.718$\pm$0.004 0.443$\pm$0.012 0.464$\pm$0.012 0.484$\pm$0.014 0.380$\pm$0.011 0.384$\pm$0.013 0.340$\pm$0.017 0.335$\pm$0.011 0.285$\pm$0.015 0.301$\pm$0.006
SUGRL 0.693$\pm$0.002 0.439$\pm$0.010 0.467$\pm$0.010 0.480$\pm$0.012 0.365$\pm$0.013 0.385$\pm$0.011 0.341$\pm$0.009 0.327$\pm$0.011 0.275$\pm$0.011 0.295$\pm$0.011
MERIT 0.717$\pm$0.004 0.442$\pm$0.009 0.463$\pm$0.009 0.483$\pm$0.010 0.368$\pm$0.011 0.381$\pm$0.011 0.341$\pm$0.012 0.324$\pm$0.012 0.272$\pm$0.010 0.304$\pm$0.009
ARIEL 0.717$\pm$0.004 0.448$\pm$0.013 0.471$\pm$0.013 0.482$\pm$0.011 0.379$\pm$0.014 0.384$\pm$0.015 0.342$\pm$0.015 0.334$\pm$0.014 0.280$\pm$0.013 0.300$\pm$0.010
SFA 0.718$\pm$0.009 0.445$\pm$0.012 0.463$\pm$0.013 0.486$\pm$0.012 0.368$\pm$0.011 0.378$\pm$0.014 0.338$\pm$0.015 0.325$\pm$0.014 0.273$\pm$0.012 0.302$\pm$0.013
Sel-Cl 0.719$\pm$0.002 0.447$\pm$0.007 0.469$\pm$0.007 0.486$\pm$0.010 0.375$\pm$0.008 0.389$\pm$0.025 0.344$\pm$0.013 0.331$\pm$0.008 0.284$\pm$0.019 0.304$\pm$0.012
Jo-SRC 0.715$\pm$0.005 0.445$\pm$0.011 0.466$\pm$0.009 0.481$\pm$0.010 0.377$\pm$0.013 0.387$\pm$0.013 0.340$\pm$0.013 0.333$\pm$0.013 0.282$\pm$0.018 0.297$\pm$0.009
GRAND+ 0.725$\pm$0.004 0.445$\pm$0.008 0.466$\pm$0.011 0.481$\pm$0.011 0.378$\pm$0.010 0.385$\pm$0.012 0.344$\pm$0.010 0.332$\pm$0.010 0.282$\pm$0.016 0.303$\pm$0.009
LR-GCL 0.728$\pm$0.006 0.472$\pm$0.013 0.492$\pm$0.011 0.508$\pm$0.014 0.405$\pm$0.014 0.411$\pm$0.012 0.405$\pm$0.012 0.359$\pm$0.015 0.307$\pm$0.013 0.335$\pm$0.013
5.1 Experimental Settings
In our experiment, we adopt eight widely used graph benchmark datasets, namely Cora, Citeseer, PubMed [Sen et al., 2008], Coauthor CS, ogbn-arxiv [Hu et al., 2020], Wiki-CS [Mernyei and Cangea, 2020], Amazon-Computers, and Amazon-Photos [Shchur et al., 2018] for the evaluation in node classification. Details of the datasets are deferred in Section 5.2 of the supplementary. Due to the fact that all public benchmark graph datasets do not come with corrupted labels or attribute noise, we manually inject noise into public datasets to evaluate our algorithm. We follow the commonly used label noise generation methods from the existing work [Han et al., 2020, Dai et al., 2022, Qian et al., 2022] to inject label noise. We generate noisy labels over all classes in two types: (1) Symmetric, where nodes from each class is flipped to other classes with a uniform random probability; (2) Asymmetric, where mislabeling only occurs between similar classes. The percentage of nodes with flipped labels is defined as the label noise level in our experiments. To evaluate the performance of our method with attribute noise, we randomly shuffle a certain percentage of input attributes for each node following [Ding et al., 2022]. The percentage of shuffled attributes is defined as the attribute noise level in our experiments.
Table 2: Ablation study on the value of rank $r_{0}$ in the optimization problem (3) on Cora with different levels of asymmetric and symmetric label noise. The accuracy with the optimal rank is shown in the last row. The accuracy difference against the optimal rank is shown for other ranks.
Rank Noise Level
0 40 60 80
- Asymmetric Symmetric Asymmetric Symmetric Asymmetric Symmetric
0.1 $\min\left\{N,d\right\}$0.002 -0.001 -0.002 -0.002 -0.001 -0.001 -0.000
0.2 $\min\left\{N,d\right\}$$0$$0$$0$$0$$0$$0$0.000
0.3 $\min\left\{N,d\right\}$$0$$0$0.001 -0.002 -0.001 -$0$0.001
0.4 $\min\left\{N,d\right\}$0.001 -0.003 -0.002 -0.001 -0.002 -0.002 -0.002
0.5 $\min\left\{N,d\right\}$0.001 -0.002 -0.003 -0.003 -0.003 -0.001 -0.002
0.6 $\min\left\{N,d\right\}$0.003 -0.002 -0.002 -0.003 -0.002 -0.002 -0.003
0.7 $\min\left\{N,d\right\}$0.003 -0.004 -0.003 -0.004 -0.004 -0.004 -0.005
0.8 $\min\left\{N,d\right\}$0.002 -0.005 -0.006 -0.006 -0.006 -0.007 -0.007
0.9 $\min\left\{N,d\right\}$0.004 -0.004 -0.005 -0.007 -0.008 -0.008 -0.006
$\min\left\{N,d\right\}$0.004 -0.004 -0.007 -0.007 -0.008 -0.010 -0.008
optimal 0.858 0.589 0.713 0.492 0.587 0.306 0.419
Tuning $r_{0},\tau$ by Cross-Validation. We tune the rank $r_{0}$ and the weight for the truncated nuclear loss $\tau$ by standard cross-validation on each dataset. Let $r_{0}=\lceil\gamma\min\left\{N,d\right\}\rceil$ where $\gamma$ is the rank ratio. We select the values of $\gamma$ and $\tau$ by performing $5$-fold cross-validation on $20\%$ of the training data in each dataset. The value of $\gamma$ is selected from $\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9\}$. The value of $\tau$ is selected from $\{0.05,0.1,0.15,0.2,0.25,0.3,0.35,0.4,0.45,0.5\}$. The selected values on each dataset are shown in Table 3.
Table 3: Selected rank ratio $\gamma$ and truncated nuclear loss’s weight $\lambda$ for each dataset.
Hyper-parameters Cora Citeseer PubMed Coauthor CS ogbn-arxiv Wiki-CS Amazon-Computers Amazon-Photos
$\tau$ 0.10 0.10 0.10 0.20 0.10 0.25 0.20 0.20
$\gamma$ 0.2 0.2 0.3 0.3 0.4 0.2 0.2 0.3
Table 4: The statistics of the datasets.
Dataset Nodes Edges Features Classes
Cora 2,708 5,429 1,433 7
CiteSeer 3,327 4,732 3,703 6
PubMed 19,717 44,338 500 3
Coauthor CS 18,333 81,894 6,805 15
ogbn-arxiv 169,343 1,166,243 128 40
Wiki-CS 11,701 215,863 300 10
Amazon-Computers 13,752 245,861 767 10
Amazon-Photos 7,650 119,081 745 8
5.2 Datasets
We evaluate our method on eight public benchmarks that are widely used for node representation learning, namely Cora, Citeseer, PubMed [Sen et al., 2008], Coauthor CS, ogbn-arxiv [Hu et al., 2020], Wiki-CS [Mernyei and Cangea, 2020], Amazon-Computers, and Amazon-Photos [Shchur et al., 2018]. Cora, Citeseer, and PubMed are the three most widely used citation networks. Coauthor CS is a co-authorship graph. The ogbn-arxiv is a directed citation graph. Wiki-CS is a hyperlink networks of computer science articles. Amazon-Computers and Amazon-Photos are co-purchase networks of products selling on Amazon.com. We summarize the statistics of all the datasets in Table 4. For all our experiments, we follow the default separation [Shchur et al., 2018, Mernyei and Cangea, 2020, Hu et al., 2020] of training, validation, and test sets on each benchmark.
5.3 Node Classification
[…]
Experimental Results.
We first compare LR-GCL against competing methods for semi-supervised or transductive node classification on input with two types of label noise. To show the robustness of LR-GCL against label noise, we perform the experiments on graphs injected with different levels of label noise ranging from $40\%$ to $80\%$ with a step of $20\%$. We follow the widely used semi-supervised setting [Kipf and Welling, 2017] for node classification. In LR-GCL, we train a transductive classifier for node classification. Previous GCL methods, including MERIT, SUGRL, and SFA, train a linear layer for inductive classification on top of the node representations learned by contrastive learning without using test data in training. Because LR-GCL is a transductive classifier, for fair comparisons, we also train the compared GCL baselines with the same transductive classifier as that for LR-GCL and a two-layer GCN transductive classifier. The results with different types of classifiers are deferred in Section C.2 of the supplementary. For all the baselines in our experiments which perform inductive classification when predicting tbe labels, we report their best results among using their original inductive classifier and two types of transductive classifiers: the same transductive classifier as that for LR-GCL and a two-layer GCN transductive classifier.
Results on Cora, Citeseer, PubMed, Coauthor-CS, and ogbn-arxiv are shown in Table 1, where we report the means of the accuracy of $10$ runs and the standard deviation. Results on Wiki-CS, Amazon-Computers, and Amazon-Photos are deferred in Section C.1 of the supplementary. It is observed from the results that LR-GCL outperforms all the baselines. By selecting confident nodes and computing robust prototypes using BEC, LR-GCL outperforms all the baselines by an even larger margin with a larger label noise level. In addition, we compare LR-GCL with baselines for noisy input with attribute noise levels ranging from $40\%$ to $80\%$ with a step of $20\%$. Results on Cora, Citeseer, and Coauthor CS are shown in Table 5 in the supplementary, where we report the means of the accuracy of $10$ runs and the standard deviation. The results clearly show that LR-GCL is more robust to attribute noise compared to all the baselines for different noise levels.
Additional Results and Ablation Studies We compare the training time of LR-GCL with competing baselines in Table 7 of the supplementary. We also perform ablation study on the value of rank $r_{0}$ in the truncated nuclear norm ${\left\|\mathbf{H}\right\|}_{r_{0}}$ in the loss function (2) of LR-GCL. It is observed from Table 2 that the performance of our LR-GCL is consistently close to the best performance among all the choices of the rank when $r_{0}$ is between $0.1\min\left\{N,d\right\}$ and $0.3\min\left\{N,d\right\}$.
6 Conclusions
In this paper, we propose a novel GCL encoder termed Low-Rank Graph Contrastive Learning (LR-GCL). LR-GCL is a robust GCL encoder which produces low-rank features inspired by the low frequency property of universal graph datasets and the sharp generalization bound for transductive learning. LR-GCL is trained with prototypical GCL with the truncated nuclear norm as the regularization term. We evaluate the performance of LR-GCL with comparison to competing baselines on semi-supervised or transductive node classification, where graph data are corrupted with noise in either the labels for the node attributes. Extensive experimental results demonstrate that LR-GCL generates more robust node representations with better performance than the current state-of-the-art node representation learning methods.
Appendix A Theoretical Results
First, recall that the optimization problem of the low-rank transductive classification is
$\min\limits_{\mathbf{W}}\frac{1}{m}\sum_{v_{i}\in\mathcal{V_{L}}}\textup{KL}\left(\mathbf{y}_{i},\left[\textup{softmax}\left(\mathbf{H}\mathbf{W}\right)\right]_{i}\right).$
We then present the proof of Theorem 4.1.
Proof of Theorem 4.1.
It can be verified that at the $t$-th iteration of gradient descent for $t\geq 1$, we have
$\mathbf{W}^{(t)}=\mathbf{W}^{(t-1)}-\eta\mathbf{H}^{\top}\left(\mathbf{H}\mathbf{W}^{(t-1)}-\mathbf{Y}\right).$ (5)
It follows by (5) that
$\mathbf{H}\mathbf{W}^{(t)}=\mathbf{H}\mathbf{W}^{(t-1)}-\eta\mathbf{K}\left(\mathbf{H}\mathbf{W}^{(t-1)}-\mathbf{Y}\right),$ (6)
where $\mathbf{K}=\mathbf{H}\mathbf{H}^{\top}$. With $\mathbf{F}(\mathbf{W},t)=\mathbf{H}\mathbf{W}^{(t)}$, it follows by (6) that
$\mathbf{F}(\mathbf{W},t)-\mathbf{Y}=\left(\mathbf{I}_{N}-\eta\mathbf{K}\right)\left(\mathbf{F}(\mathbf{W},r,t-1)-\mathbf{Y}\right).$
It follows from the above equality that
$\mathbf{F}(\mathbf{W},t)-\mathbf{Y}=\left(\mathbf{I}_{N}-\eta\mathbf{K}\right)^{t}\left(\mathbf{F}(\mathbf{W},0)-\mathbf{Y}\right)=-\left(\mathbf{I}_{N}-\eta\mathbf{K}\right)^{t}\mathbf{Y},$
so that
${\left\|\mathbf{F}(\mathbf{W},t)-\mathbf{Y}\right\|}_{\textup{F}}\leq{\left\|\left(\mathbf{I}_{N}-\eta\mathbf{K}\right)^{t}\mathbf{Y}\right\|}_{\textup{F}}.$ (7)
As a result of (7), we have
${\left\|\left[\mathbf{F}(\mathbf{W},t)-\mathbf{Y}\right]_{{\cal L}}\right\|}_{\textup{F}}\leq{\left\|\left(\mathbf{I}_{N}-\eta\mathbf{K}\right)^{t}\mathbf{Y}\right\|}_{\textup{F}}.$ (8)
We apply [Yang, 2023, Corollary 3.7] to obtain the following bound for the test loss $\frac{1}{u}{\left\|\left[\mathbf{F}(\mathbf{W},t)-\mathbf{Y}\right]_{{\cal U}}\right\|}_{\textup{F}}^{2}$:
$\frac{1}{u}{\left\|\left[\mathbf{F}(\mathbf{W},t)-\mathbf{Y}\right]_{{\cal U}}\right\|}_{\textup{F}}^{2}$ $\leq\frac{c_{0}}{m}{\left\|\left[\mathbf{F}(\mathbf{W},t)-\mathbf{Y}\right]_{{\cal L}}\right\|}_{\textup{F}}^{2}+c_{1}\min_{0\leq Q\leq n}r(u,m,Q)+\frac{c_{2}x}{u},$ (9)
with
$r(u,m,Q)\coloneqq Q\left(\frac{1}{u}+\frac{1}{m}\right)+\left(\sqrt{\frac{\sum\limits_{q=Q+1}^{n}\widehat{\lambda}_{q}}{u}}+\sqrt{\frac{\sum\limits_{q=Q+1}^{n}\widehat{\lambda}_{q}}{m}}\right).$
where $c_{0},c_{1},c_{2}$ are positive numbers depending on $\mathbf{U}$, $\left\{\widehat{\lambda}_{i}\right\}_{i=1}^{r}$, and $\tau_{0}$ with $\tau_{0}^{2}=\max_{i\in[N]}\mathbf{K}_{ii}$.
It follows by (7) and (9) with $Q=r_{0}$ that
$\frac{1}{u}{\left\|\left[\mathbf{F}(\mathbf{W},t)-\mathbf{Y}\right]_{{\cal U}}\right\|}_{\textup{F}}^{2}$
$\leq\frac{c_{0}}{m}{\left\|\left[\mathbf{F}(\mathbf{W},t)-\mathbf{Y}\right]_{{\cal L}}\right\|}_{\textup{F}}^{2}+c_{1}r_{0}\left(\frac{1}{u}+\frac{1}{m}\right)+\left(\sqrt{\frac{\sum\limits_{q=r_{0}+1}^{n}\widehat{\lambda}_{q}}{u}}+\sqrt{\frac{\sum\limits_{q=r_{0}+1}^{n}\widehat{\lambda}_{q}}{m}}\right)+\frac{c_{2}x}{u},$ (10)
which completes the proof.
Appendix B Implementation Details
B.1 Compared Methods
To demonstrate the power of LR-GCL in learning robust node representation, we compare LR-GCL with two robust contrastive learning baselines, Jo-SRC and Sel-CL, which select clean samples for image data. Since their sample selection methods are general and not limited to the image domain, we adopt these two baselines in our experiments. Next, we introduce the implementation details of applying Jo-SRC and Sel-CL on graph data.
Jo-SRC [Yao et al., 2021]: Jo-SRC is a robust contrastive learning method proposed for image classification. It selects clean samples for training by adopting the Jensen-Shannon divergence to measure the likelihood of each sample being clean. Because this method is a general selection strategy on the representation space, it is adapted to selecting clean samples in the representation space of nodes in GCL. It also introduces a consistency regularization term to the contrastive loss to improve the robustness. To get a competitive robust GCL baseline, we apply the sample selection strategy and the consistency regularization proposed by Jo-SRC to state-of-the-art GCL methods MVGRL, MERIT, and SUGRL. We add the regularization term in Jo-SRC to the graph contrastive loss. The GCL encoders are trained only on the clean samples selected by Jo-SRC. We only report the best results for comparison, which are achieved by applying Jo-SRC to MERIT.
Sel-CL [Li et al., 2022]: Sel-CL is a supervised contrastive learning proposed to learn robust pre-trained representations for image classification. It proposes to select confident contrastive pairs in the contrastive learning frameworks. Sel-CL first selects confident examples by measuring the agreement between learned representations and labels generated by label propagation with cross-entropy loss. Next, Sel-CL selects contrastive pairs from those with selected confident examples in them. This method is also a general sample selection strategy on a learned representation space. So we can adapt Sel-CL to the node representation space to select confident pairs for GCL. In this process, they only select contrastive pairs whose representation similarity is higher than a dynamic threshold. In our experiments, we also adopt the confident contrastive pair selection strategy to the state-of-the-art GCL methods MVGRL, MERIT, and SUGRL. With the same GCL framework, GCL encoders are trained only on the confident pairs selected by Sel-CL. We only report the best results for comparison, which are achieved by applying Sel-CL to MERIT.
Appendix C Additional Experimental Results
C.1 Node Classification on Wiki-CS, Amazon-Computers, and Amazon-Photos
The results for node classification with symmetric label noise, asymmetric label noise, and attribute noise on Wiki-CS, Amazon-Computers, and Amazon-Photos are shown in Table 5. It is observed that LR-GCL also outperforms all the baselines for node classification with both label noise and attribute noise on these three benchmark datasets.
Table 5: Performance comparison for node classification on Wiki-CS, Amazon-Computers, and Amazon-Photos with asymmetric label noise, symmetric label noise, and attribute noise.
Noise Level
0 40 60 80
Dataset Methods - Asymmetric Symmetric Attribute Asymmetric Symmetric Attribute Asymmetric Symmetric Attribute
Wiki-CS GCN 0.918$\pm$0.001 0.645$\pm$0.009 0.656$\pm$0.006 0.702$\pm$0.010 0.511$\pm$0.013 0.501$\pm$0.009 0.531$\pm$0.010 0.429$\pm$0.022 0.389$\pm$0.011 0.415$\pm$0.013
[…]
GCE 0.922$\pm$0.003 0.662$\pm$0.017 0.659$\pm$0.007 0.705$\pm$0.014 0.515$\pm$0.016 0.502$\pm$0.007 0.539$\pm$0.009 0.443$\pm$0.017 0.389$\pm$0.012 0.412$\pm$0.011
UnionNET 0.918$\pm$0.002 0.669$\pm$0.023 0.671$\pm$0.013 0.706$\pm$0.012 0.525$\pm$0.011 0.529$\pm$0.011 0.540$\pm$0.012 0.458$\pm$0.015 0.401$\pm$0.011 0.420$\pm$0.007
NRGNN 0.919$\pm$0.002 0.678$\pm$0.014 0.689$\pm$0.009 0.705$\pm$0.012 0.545$\pm$0.021 0.556$\pm$0.011 0.546$\pm$0.011 0.461$\pm$0.012 0.410$\pm$0.012 0.417$\pm$0.007
RTGNN 0.920$\pm$0.005 0.678$\pm$0.012 0.691$\pm$0.009 0.712$\pm$0.008 0.559$\pm$0.010 0.569$\pm$0.011 0.560$\pm$0.008 0.455$\pm$0.015 0.415$\pm$0.015 0.412$\pm$0.014
SUGRL 0.922$\pm$0.005 0.675$\pm$0.010 0.695$\pm$0.010 0.714$\pm$0.006 0.550$\pm$0.011 0.560$\pm$0.011 0.561$\pm$0.007 0.449$\pm$0.011 0.411$\pm$0.011 0.429$\pm$0.008
MERIT 0.924$\pm$0.004 0.679$\pm$0.011 0.689$\pm$0.008 0.709$\pm$0.005 0.552$\pm$0.014 0.562$\pm$0.014 0.562$\pm$0.011 0.452$\pm$0.013 0.403$\pm$0.013 0.426$\pm$0.005
ARIEL 0.925$\pm$0.004 0.682$\pm$0.011 0.699$\pm$0.009 0.712$\pm$0.005 0.555$\pm$0.011 0.566$\pm$0.011 0.556$\pm$0.011 0.454$\pm$0.014 0.415$\pm$0.019 0.427$\pm$0.013
SFA 0.925$\pm$0.009 0.682$\pm$0.011 0.690$\pm$0.012 0.715$\pm$0.012 0.555$\pm$0.015 0.567$\pm$0.014 0.565$\pm$0.013 0.458$\pm$0.013 0.402$\pm$0.013 0.429$\pm$0.015
Sel-Cl 0.922$\pm$0.008 0.684$\pm$0.009 0.694$\pm$0.012 0.714$\pm$0.010 0.557$\pm$0.013 0.568$\pm$0.013 0.566$\pm$0.010 0.457$\pm$0.013 0.412$\pm$0.017 0.425$\pm$0.009
Jo-SRC 0.921$\pm$0.005 0.684$\pm$0.011 0.695$\pm$0.004 0.709$\pm$0.007 0.560$\pm$0.011 0.566$\pm$0.011 0.561$\pm$0.009 0.456$\pm$0.013 0.410$\pm$0.018 0.428$\pm$0.010
GRAND+ 0.927$\pm$0.004 0.682$\pm$0.011 0.693$\pm$0.006 0.715$\pm$0.008 0.554$\pm$0.008 0.568$\pm$0.013 0.557$\pm$0.011 0.455$\pm$0.012 0.416$\pm$0.013 0.428$\pm$0.011
LR-GCL 0.933$\pm$0.006 0.699$\pm$0.015 0.721$\pm$0.011 0.742$\pm$0.015 0.575$\pm$0.014 0.595$\pm$0.018 0.588$\pm$0.015 0.469$\pm$0.015 0.438$\pm$0.015 0.453$\pm$0.017
Amazon-Computers GCN 0.815$\pm$0.005 0.547$\pm$0.015 0.636$\pm$0.007 0.639$\pm$0.008 0.405$\pm$0.014 0.517$\pm$0.010 0.439$\pm$0.012 0.265$\pm$0.012 0.354$\pm$0.014 0.317$\pm$0.013
[…]
GCE 0.819$\pm$0.004 0.573$\pm$0.011 0.652$\pm$0.008 0.650$\pm$0.014 0.449$\pm$0.011 0.509$\pm$0.011 0.445$\pm$0.015 0.280$\pm$0.013 0.353$\pm$0.013 0.325$\pm$0.015
UnionNET 0.820$\pm$0.006 0.569$\pm$0.014 0.664$\pm$0.007 0.653$\pm$0.012 0.452$\pm$0.010 0.541$\pm$0.010 0.450$\pm$0.009 0.283$\pm$0.014 0.370$\pm$0.011 0.320$\pm$0.012
NRGNN 0.822$\pm$0.006 0.571$\pm$0.019 0.676$\pm$0.007 0.645$\pm$0.012 0.470$\pm$0.014 0.548$\pm$0.014 0.451$\pm$0.011 0.282$\pm$0.022 0.373$\pm$0.012 0.326$\pm$0.010
RTGNN 0.828$\pm$0.003 0.570$\pm$0.010 0.682$\pm$0.008 0.678$\pm$0.011 0.474$\pm$0.011 0.555$\pm$0.010 0.457$\pm$0.009 0.280$\pm$0.011 0.386$\pm$0.014 0.342$\pm$0.016
SUGRL 0.834$\pm$0.005 0.564$\pm$0.011 0.674$\pm$0.012 0.675$\pm$0.009 0.468$\pm$0.011 0.552$\pm$0.011 0.452$\pm$0.012 0.280$\pm$0.012 0.381$\pm$0.012 0.338$\pm$0.014
MERIT 0.831$\pm$0.005 0.560$\pm$0.008 0.670$\pm$0.008 0.671$\pm$0.009 0.467$\pm$0.013 0.547$\pm$0.013 0.450$\pm$0.014 0.277$\pm$0.013 0.385$\pm$0.013 0.335$\pm$0.009
ARIEL 0.843$\pm$0.004 0.573$\pm$0.013 0.681$\pm$0.010 0.675$\pm$0.009 0.471$\pm$0.012 0.553$\pm$0.012 0.455$\pm$0.014 0.284$\pm$0.014 0.389$\pm$0.013 0.343$\pm$0.013
SFA 0.839$\pm$0.010 0.564$\pm$0.011 0.677$\pm$0.013 0.676$\pm$0.015 0.473$\pm$0.014 0.549$\pm$0.014 0.457$\pm$0.014 0.282$\pm$0.016 0.389$\pm$0.013 0.344$\pm$0.017
Sel-Cl 0.828$\pm$0.002 0.570$\pm$0.010 0.685$\pm$0.012 0.676$\pm$0.009 0.472$\pm$0.013 0.554$\pm$0.014 0.455$\pm$0.011 0.282$\pm$0.017 0.389$\pm$0.013 0.341$\pm$0.015
Jo-SRC 0.825$\pm$0.005 0.571$\pm$0.006 0.684$\pm$0.013 0.679$\pm$0.007 0.473$\pm$0.011 0.556$\pm$0.008 0.458$\pm$0.012 0.285$\pm$0.013 0.387$\pm$0.018 0.345$\pm$0.018
GRAND+ 0.858$\pm$0.006 0.570$\pm$0.009 0.682$\pm$0.007 0.678$\pm$0.011 0.472$\pm$0.010 0.554$\pm$0.008 0.456$\pm$0.012 0.284$\pm$0.015 0.387$\pm$0.015 0.345$\pm$0.013
LR-GCL 0.858$\pm$0.006 0.589$\pm$0.011 0.713$\pm$0.007 0.695$\pm$0.011 0.492$\pm$0.011 0.587$\pm$0.013 0.477$\pm$0.012 0.306$\pm$0.012 0.419$\pm$0.012 0.363$\pm$0.011
Amazon-Photos GCN 0.703$\pm$0.005 0.475$\pm$0.023 0.501$\pm$0.013 0.529$\pm$0.009 0.351$\pm$0.014 0.341$\pm$0.014 0.372$\pm$0.011 0.291$\pm$0.022 0.281$\pm$0.019 0.290$\pm$0.014
[…]
GCE 0.705$\pm$0.004 0.490$\pm$0.016 0.512$\pm$0.014 0.540$\pm$0.014 0.362$\pm$0.015 0.352$\pm$0.010 0.381$\pm$0.009 0.309$\pm$0.012 0.285$\pm$0.014 0.285$\pm$0.011
UnionNET 0.706$\pm$0.006 0.499$\pm$0.015 0.547$\pm$0.014 0.545$\pm$0.013 0.379$\pm$0.013 0.399$\pm$0.013 0.379$\pm$0.012 0.322$\pm$0.021 0.302$\pm$0.013 0.290$\pm$0.012
NRGNN 0.710$\pm$0.006 0.498$\pm$0.015 0.546$\pm$0.015 0.538$\pm$0.011 0.382$\pm$0.016 0.412$\pm$0.016 0.377$\pm$0.012 0.336$\pm$0.021 0.309$\pm$0.018 0.284$\pm$0.009
RTGNN 0.746$\pm$0.008 0.498$\pm$0.007 0.556$\pm$0.007 0.550$\pm$0.012 0.392$\pm$0.010 0.424$\pm$0.013 0.390$\pm$0.014 0.348$\pm$0.017 0.308$\pm$0.016 0.302$\pm$0.011
SUGRL 0.730$\pm$0.005 0.493$\pm$0.011 0.541$\pm$0.011 0.544$\pm$0.010 0.376$\pm$0.009 0.421$\pm$0.009 0.388$\pm$0.009 0.339$\pm$0.010 0.305$\pm$0.010 0.300$\pm$0.009
MERIT 0.740$\pm$0.007 0.496$\pm$0.012 0.536$\pm$0.012 0.542$\pm$0.010 0.383$\pm$0.011 0.425$\pm$0.011 0.387$\pm$0.008 0.344$\pm$0.014 0.301$\pm$0.014 0.295$\pm$0.009
SFA 0.740$\pm$0.011 0.502$\pm$0.014 0.532$\pm$0.015 0.547$\pm$0.013 0.390$\pm$0.014 0.433$\pm$0.014 0.389$\pm$0.012 0.347$\pm$0.016 0.312$\pm$0.015 0.299$\pm$0.013
ARIEL 0.727$\pm$0.007 0.500$\pm$0.008 0.550$\pm$0.013 0.548$\pm$0.008 0.391$\pm$0.009 0.427$\pm$0.012 0.389$\pm$0.014 0.349$\pm$0.014 0.307$\pm$0.013 0.299$\pm$0.013
Sel-Cl 0.725$\pm$0.008 0.499$\pm$0.012 0.551$\pm$0.010 0.549$\pm$0.008 0.389$\pm$0.011 0.426$\pm$0.008 0.391$\pm$0.020 0.350$\pm$0.018 0.310$\pm$0.015 0.300$\pm$0.017
Jo-SRC 0.730$\pm$0.005 0.500$\pm$0.013 0.555$\pm$0.011 0.551$\pm$0.011 0.394$\pm$0.013 0.425$\pm$0.013 0.393$\pm$0.013 0.351$\pm$0.013 0.305$\pm$0.018 0.303$\pm$0.013
GRAND+ 0.756$\pm$0.004 0.497$\pm$0.010 0.553$\pm$0.010 0.552$\pm$0.011 0.390$\pm$0.013 0.422$\pm$0.013 0.387$\pm$0.013 0.348$\pm$0.013 0.309$\pm$0.014 0.302$\pm$0.012
LR-GCL 0.757$\pm$0.010 0.520$\pm$0.013 0.581$\pm$0.013 0.570$\pm$0.007 0.410$\pm$0.014 0.455$\pm$0.014 0.406$\pm$0.012 0.369$\pm$0.012 0.335$\pm$0.014 0.318$\pm$0.010
C.2 Node Classification Results for GCL Methods with different types of classifiers
Existing GCL methods, such as MERIT, SUGRL, and SFA, first train a graph encoder with graph contrastive learning objectives such as InfoNCE [Jin et al., 2021]. After obtaining the node representation learned by contrastive learning, a linear layer for classification is trained in the supervised setting. In contrast, LR-GCL adopts a transductive classifier on top of the node representation obtained by contrastive learning. For fair comparisons with previous GCL methods, we also train the compared GCL baselines with the same transductive classifier as in LR-GCL and a two-layer transductive GCN classifier. The results with different types of classifiers are deferred in Section C.2 of the supplementary.
Table 6: Performance comparison for node classification by inductive linear classifier, transductive two-layer GCN classifier, and transductive classifier used in LR-GCL. The comparisons are performed on Cora with asymmetric label noise, symmetric label noise, and attribute noise.
Noise Level
0 40 60 80
Methods - Asymmetric Symmetric Attribute Asymmetric Symmetric Attribute Asymmetric Symmetric Attribute
SUGRL (original, inductive classifier) 0.834$\pm$0.005 0.564$\pm$0.011 0.674$\pm$0.012 0.675$\pm$0.009 0.468$\pm$0.011 0.552$\pm$0.011 0.452$\pm$0.012 0.280$\pm$0.012 0.381$\pm$0.012 0.338$\pm$0.014
SUGRL + transductive GCN 0.833$\pm$0.006 0.562$\pm$0.013 0.675$\pm$0.015 0.673$\pm$0.012 0.470$\pm$0.011 0.551$\pm$0.011 0.454$\pm$0.012 0.280$\pm$0.012 0.380$\pm$0.012 0.340$\pm$0.014
SUGRL + linear transductive classifier 0.836$\pm$0.007 0.568$\pm$0.013 0.677$\pm$0.010 0.674$\pm$0.011 0.472$\pm$0.011 0.555$\pm$0.011 0.457$\pm$0.012 0.284$\pm$0.012 0.383$\pm$0.012 0.341$\pm$0.014
MERIT (original, inductive classifier) 0.831$\pm$0.005 0.560$\pm$0.008 0.670$\pm$0.008 0.671$\pm$0.009 0.467$\pm$0.013 0.547$\pm$0.013 0.450$\pm$0.014 0.277$\pm$0.013 0.385$\pm$0.013 0.335$\pm$0.009
MERIT + transductive GCN 0.831$\pm$0.007 0.562$\pm$0.011 0.668$\pm$0.013 0.672$\pm$0.014 0.466$\pm$0.013 0.549$\pm$0.015 0.451$\pm$0.016 0.276$\pm$0.012 0.382$\pm$0.014 0.337$\pm$0.013
MERIT + linear transductive classifier 0.833$\pm$0.003 0.562$\pm$0.014 0.673$\pm$0.012 0.673$\pm$0.011 0.466$\pm$0.015 0.546$\pm$0.016 0.453$\pm$0.017 0.280$\pm$0.016 0.386$\pm$0.011 0.336$\pm$0.014
SFA (original, inductive classifier) 0.839$\pm$0.010 0.564$\pm$0.011 0.677$\pm$0.013 0.676$\pm$0.015 0.473$\pm$0.014 0.549$\pm$0.014 0.457$\pm$0.014 0.282$\pm$0.016 0.389$\pm$0.013 0.344$\pm$0.017
SFA + transductive GCN 0.837$\pm$0.013 0.565$\pm$0.011 0.673$\pm$0.017 0.673$\pm$0.018 0.474$\pm$0.016 0.551$\pm$0.015 0.453$\pm$0.018 0.277$\pm$0.016 0.389$\pm$0.015 0.343$\pm$0.019
SFA + linear transductive classifier 0.841$\pm$0.015 0.566$\pm$0.013 0.678$\pm$0.014 0.679$\pm$0.014 0.477$\pm$0.015 0.552$\pm$0.012 0.456$\pm$0.016 0.284$\pm$0.017 0.391$\pm$0.015 0.348$\pm$0.019
LR-GCL 0.858$\pm$0.006 0.589$\pm$0.011 0.713$\pm$0.007 0.695$\pm$0.011 0.492$\pm$0.011 0.587$\pm$0.013 0.477$\pm$0.012 0.306$\pm$0.012 0.419$\pm$0.012 0.363$\pm$0.011
C.3 Training Time Comparisons and Study on Different Ranks
In this section, we first compare the training time of LR-GCL against other baseline methods on all benchmark datasets. For our method, we include the training time of robust graph contrastive learning, the time of the SVD computation of the kernel, and the training time of the transductive classifier. For graph contrastive learning methods, we include both the training time of the GCL encoder and the downstream classifier. We evaluate the training time on one 80 GB A100 GPU. The results are shown in Table 7. It is observed that the LR-GCL takes similar training time as baseline GCL methods such as SFA and MERIT.
We also perform ablation study on the value of rank $r$ in the optimization problem (3) for our low-rank transductive classifier. It is observed from Table 2 that the performance of our low-rank classifier is consistently close to the best performance among all the choices of the rank when $r$ is between $0.1\min\left\{N,d\right\}$ and $0.2\min\left\{N,d\right\}$.
Table 7: Training time (seconds) comparisons for node classification.
Methods Cora Citeseer PubMed Coauthor-CS Wiki-CS Computer Photo ogbn-arxiv
GCN 11.5 13.7 38.6 43.2 22.3 30.2 19.0 215.1
[…]
GCE 32.6 36.9 67.3 80.8 37.6 50.1 32.2 346.1
UnionNET 67.5 69.7 100.5 124.2 53.2 69.2 45.3 479.3
NRGNN 72.4 80.5 142.7 189.4 74.3 97.2 62.4 650.2
RTGNN 143.3 169.5 299.5 353.5 153.7 201.5 124.2 1322.2
SUGRL 100.3 122.1 207.4 227.1 107.7 142.8 87.7 946.8
MERIT 167.2 179.2 336.7 375.3 172.3 226.5 140.6 1495.1
ARIEL 156.9 164.3 284.3 332.6 145.1 190.4 118.3 1261.4
SFA 237.5 269.4 457.1 492.3 233.5 304.5 187.2 2013.1
Sel-Cl 177.3 189.9 313.5 352.5 161.7 211.1 130.9 1401.1
Jo-SRC 148.2 157.1 281.0 306.1 144.5 188.0 118.5 1256.0
GRAND+ 57.4 68.4 101.7 124.2 54.8 73.8 44.5 479.2
LR-GCL 159.9 174.5 350.7 380.9 180.3 235.7 145.5 1552.7
C.4 Eigen-Projection and Concentration Entropy
Figure 3 illustrates the eigen-projection and energy concentration ratio for more datasets.
Figure 3: Eigen-projection (first row) and energy concentration (second row) on Coauthor-CS, Amazon-Computers, Amazon-Photos, and ogbn-arxiv. By the rank of $0.2\min\left\{N,d\right\}$, the concentration entropy on Coauthor-CS, Amazon-Computers, Amazon-Photos, and ogbn-arxiv are $0.779$, $0.809$, $0.752$, and $0.787$.
C.5 Visualization of Confidence Score
We visualize the confident nodes selected by BPL in the embedding space of the learned node representations in Figure 4. The node representations are visualized by the t-SNE figure. Each mark in t-SNE represents the representation of a node, and the color of the mark denotes the confidence of that node. The results are shown for different levels of attribute noise. It is observed from Figure 4 that confident nodes, which are redder in Figure 4, are well separated in the embedding space. With a higher level of attribute noise, the bluer nodes from different clusters blended around the cluster boundaries. In contrast, the redder nodes are still well separated and far away from cluster boundaries, which leads to more robustness and better performance in downstream tasks.
(a) noise level = 0
(b) noise level = 20
(c) noise level = 40
Figure 4: Visualization of confident nodes with different levels of attribute noise for semi-supervised node classification on Citeseer.
C.6 Study in the Kernel Complexity
With feature matrix $\mathbf{H}$ produced by different encoders, we compute a kernel matrix $\mathbf{K}=\widehat{}\mathbf{A}\mathbf{H}\mathbf{H}^{\top}\widehat{}\mathbf{A}^{\top}$ before and after optimization of problem (3). Score of $\mathbf{K}$ is
complexity of $\textup{complexity of }\mathbf{K}=\min_{h\in[0,n]}\frac{h}{n}+\sqrt{\frac{\sum\limits_{i=h+1}^{n}\widehat{\lambda}_{i}}{n}}$ (11)
Also, compute $\mathbf{K}$ for LR-GCL and other GCL baselines.
Datasets MERIT SFA RGCL LR-GCL
Cora Complex. 0.5485 0.5749 0.5383 0.1675
$h$ 1455 1532 1431 433 (rank)
Citeseer Complex. 0.4233 0.4546 0.4090 0.1178
$h$ 1347 1479 1330 366
PubMed Complex. 0.9770 0.1095 0.0932 0.0674
$h$ 1690 1874 1667 1183
Wiki-CS Complex. 0.1615 0.1762 0.1498 0.1062
$h$ 1784 1986 1639 1170
Amazon-Computers Complex. 0.1132 0.1217 0.1110 0.0655
$h$ 1467 1579 1433 874
Amazon-Photos Complex. 0.1875 0.2047 0.1823 0.1007
$h$ 1397 1532 1353 750
Coauthor-CS Complex. 0.0764 0.0859 0.0749 0.0622
$h$ 1657 1874 1605 1120
ogbn-arxiv Complex. 0.0124 0.0139 0.0117 0.0103
$h$ 1769 1920 1653 1354
Table 8: Comparisons in complexity of kernels.